OpenStudy (yacoub1993):

Tell whether the sequence 1/3,0,1,-2… is arithmetic, geometric, or neither. Find the next three terms of the sequence.

OpenStudy (yacoub1993):

Any help

OpenStudy (anonymous):

for a geometric series the common ration should be the same for two terms like r=n2/n1=n3/n2 r=0/(1/3) is not equal to 1/0 so it is not a geometrc series

OpenStudy (anonymous):

for a arithmetic series the common difference ashould be the same d=n2-n1=n3-n2 d=0-1/3 is not equal to 1-0 conclusion it is neither arithametic nor geometric

OpenStudy (anonymous):

did u get it ??? @Yacoub1993

OpenStudy (yacoub1993):

yup thnx buddy..

OpenStudy (yacoub1993):

sambhav__jain But these r the choices they gave me A. neither; 7, -20, 61 B. geometric;7, -20, 61 C. arithmetic; -1/3,1 1/3,3 D. geometric;-3 1/3,-5 5/9,-9 7/27

OpenStudy (yacoub1993):

@sambhav__jain